Find a function f such that 5+ dt 4.x-7 – 22
Answer #1


5+41 0== F (1) dt = 4x-?- x? (6) -dt = 4x-7-x-5 Differentiating with respect x, we get a -[4x² – x²-5] ET LA (2) - Sea) (a) =